An island has no currency; it instead has the following exchange rate:
50 bananas = 20 coconuts
30 coconuts = 12 fish
100 fish = 1 hammock
How many bananas equal 1 hammock?


Answer by josmiceli(19441)

1 banana = 2/5 of a coconut
2/5 of a coconut = (2/5)*(12/30) of a fish
= 24/150 = 12/75 of a fish
------------------------
12/75 of a fish = (12/75)*( 1/100 ) of a hammock
= 3/1875 = 1/625 of a hammock
-----------------------------
If 1 banana = 1/625 of a hammock, the
625 bananas = 1 hammock
-----------------------
-----------------------
check:
625 bananas = (2/5)*625 = 250 coconuts
250 coconuts = (2/5)*250 = 100 fish
100 fish = 1 hammock
OK